OpenStudy (anonymous):

the Length of a rectangle is 5cm more than twice the width. The perimeter of the rectangle is 34cm. Find the dimensions of the rectangle.

OpenStudy (ajprincess):

Let the length of the rectangle be l and width be w. l=2w+5 Perimeter=2(l+w) 34=2(2w+5+w) 34/2=3w+5 17=3w+5 find w and then to find l plug in the value u get in l=2w+5

OpenStudy (anonymous):

ohhhhhh okay